A drug tester claims that a drug cures a rare skin disease 77?% of the time. The claim is checked by testing the drug on 100 patients. If at least 72 patients are? cured, the claim will be accepted.

Find the probability that the claim will be rejected assuming that the? manufacturer's claim is true. Use the normal distribution to approximate the binomial distribution if possible.
Answer #1

here mean number of patients cured =np=100*0.77=77

and std deviaiton =sqrT(np(1-p))=4.2083

hence P( claim will be rejected)=P(less than 72 patients cured)=P(X<72)=P(Z<(71.5-77)/4.2083)

=P(Z<-1.31)=0.0951